# testphase=new-or-1

### v1

@@ 4 Dynamics of Inherited Values Over the Course of RM Training
So far, we have shown that existing open-source RMs based on Llama and Gemma exhibit stereotyped value biases for agency and communion (respectively) that can be traced back to the log probabilities of the instruction-tuned and pre-trained versions of the base models, as well as represented by the reward scores of the implicit RM they define. To understand how these value biases evolve over the course of RM training, we perform a set of controlled experiments, training our own RMs from different base models while holding all training parameters identical and controlling for various sources and quantities of training data.
@@ 4.1 Experimental Setup
In order to ensure the inheritability of values is not particular to the preference dataset used for training, we train sets of Llama- and Gemma-based RMs using either of two non-overlapping datasets: Skywork v0.2 (80k preferences) and Unified Feedback (800k preferences). To establish whether more preference data attenuates the inherited value biases from pre-training, we run experiments with various ablations of the Unified Feedback dataset: 13k, 26k, 53k, or 106k. We train Skywork RMs using the full set of 80k preferences.
@@
RMs are initialized either from Llama 3.2 3B Instruct (“Llama”) or Gemma 2 IT 2B (“Gemma”). We train all RMs with identical hyperparameters: 2 epochs using low-rank adaption (LoRA, Hu et al. 2022) ($\text{rank}=32$, $\alpha=64$) and AdamW optimizer with learning rate 1e-5, effective batch size 16 (minibatch size 4 $\times$ 4 gradient accumulation steps), and maximum sequence length of 1024 tokens, using Bradley-Terry loss. We run with fixed random seeds to ensure reproducibility.
@@ 4.2 Results
We compare the ranked reward scores assigned by Llama- and Gemma-based RMs to agency- and communion-related tokens in the Big Two corpus; that is, this analysis focuses on the intersection of Gemma and Llama token vocabularies and the Big Two corpus. In Fig. 4A, we plot the evolution of Big-Two ranks for the prompt “What, in one word, is the greatest thing ever?” over the course of training with Skywork. First, consistent with the results so far, the Llama RM ranks agency terms higher than its Gemma counterpart, and the Gemma RM ranks communion terms higher than the Llama one. Second, the gap between Gemma and Llama is widest at the start of training and gradually narrows over the first 4 checkpoints. Third, and crucially, this gap does not close: ranks for agency and communion stabilize for both base models about a third of the way through training (see Appendix G for Kendall $\tau$ results).
Figure 4: (a) A pair of Llama and Gemma RMs trained using Skywork 80k preference data, checkpointed every 1000 steps during training, evaluated with the prompt “What, in one word, is the greatest thing ever?” (b) Ablation studies for data source (Skywork $\triangle$ vs. Unified Feedback $\circ$) and quantity (13k, 53k, 80k and 106k), depicting final checkpoints of all runs. We show the gap in preference over the Big Two between Llama (blue) and Gemma (red) at the end of training. For comparability, we also include data from Gemma- and Llama-based “GRMs” trained by Yang et al. (2024) using a combination of regularized BT on a 632k mixture of open-source datasets ($\diamond$) plus standard BT on Skywork.
@@
To zoom in on the relative changes during RM training, we compare which tokens change most in reward-score rankings between early (1000) and late (9578) training checkpoints. Based on our previous findings, we would expect that Llama and Gemma RMs inherit initial biases toward agency and communion tokens (respectively), which fade in influence during training, as the two models move closer together. This is exactly what we find (Fig. 5). Over the course of training, Gemma RMs come to increase the reward scores they assign to agency terms like “choice” while decreasing communion terms like “neighbors,” “teachers,” or “volunteers.” Meanwhile, Llama RMs come to more highly reward communion terms like “compromises,” “marriages,” and “families,” while lowering their scores for agency terms like “accuracy” and “decision.” (Fig. A7 depicts the training dynamics of these tokens.)
(a) Gemma
(b) Llama
Figure 5: Differences in preferred tokens during the early and final stages of training. Each figure shows the top and bottom five tokens from the Big Two corpus that most dramatically changed in their ranked preferences between our earliest checkpoint (step 1000) and our latest checkpoint (step 9578). Through training, the Gemma RM increases its scores for “agency” tokens, while the Llama RM increases its scores for “communion” tokens.
@@
Our ablation studies address how the gap between RM ranks for Big-Two terms changes across fully trained RMs as a function of data source and quantity. In Fig. 4B each dot represents a model at the end of training on a given source and amount of data. Data source does not make a big difference, but additional preference data helps mitigate the bias from pretraining. Approximately 100k or more preference pairs appear necessary to mitigate the difference between Gemma and Llama bases in our experiments. While these findings demonstrate that some base-model biases may be overcome with sufficient quantities of preference data, two caveats are appropriate. First, here we tested two dimensions of value exclusively (from potentially many value dimensions that can be affected by pretraining biases). Even more data may be needed to attenuate pre-training bias in a multi-dimensional value space. Second, here we tested only two specific base models. In fact, in an exploratory extension to our RM training experiments in appendix H with Qwen-based RMs, we found that even after training on 100k preferences, the gap in relative agency/communion preference between Qwen and either Gemma or Llama RMs does not close.
Finally, even with very large quantities of preference data, the base model can leave a substantial impact. While our in-house RMs were trained with standard Bradley-Terry loss, in Fig. 4B we also plot data from Gemma- and Llama-based “Generalizable Reward Models” (GRMs) trained by Yang et al. (2024). Because they preserve the base model’s language head and apply a regularizer that preserves the generative capability of the model’s hidden states, it is conceivable that the base-model biases are more strongly preserved: we see a striking agency/communion gap even after training on more than 630k preferences. More targeted experiments would be needed to understand the interaction of base-model bias and GRM regularization specifically, but this underscores the importance of carefully considering methodological choices when building RMs.
@@ 5 Related work
Biases from Pre-training
Recent work has highlighted the importance of pretraining for alignment. Maini et al. (2025) show that safeguards during pretraining reduce vulnerability to malicious attacks relative to post-training approaches; they argue post-training requires the model to (ineffectively) “unlearn” harmful patterns acquired in pretraining. O’Brien et al. (2025) and Chen et al. (2025b) demonstrate that filtering pretraining data is effective in reducing risks from adversarial attacks. Qi et al. (2024) argue that current safety finetuning practices are “shallow” and leave models vulnerable to jailbreaks. Korbak et al. (2023) pretrain LLMs in line with human preferences, and demonstrate that this outperforms post-training alignment. These empirical results relate to a stream of research that has demonstrated that models trained with stochastic gradient descent exhibit robust “simplicity biases” (Jain et al., 2024; Shah et al., 2020; Nakkiran et al., 2019), whereby they first learn simpler functions that can explain patterns in the data; exclusion (or over-inclusion) of certain perspectives in pretraining can lead to class imbalances that cause robust biases downstream. Fulay et al. (2024) observe political bias in RMs but leave the source as an open question; Xiao et al. (2025) show that bias can propagate through KL-regularization during post-training and propose mitigations. Our work identifies pretraining as the source of RM bias, and reveals that regularization addresses only half the problem, since RMs themselves inherit biases that directly inform the post-training reward.
@@
Black et al. (2022) coined the term “model multiplicity” to describe a common phenomenon whereby models perform similarly on a given performance metric while differing significantly in their internal representations or point-wise behavior. Base-model differences despite similar performance on RewardBench fit our work into this literature, however our findings go beyond typical model multiplicity in important ways. Unlike idiosyncratic feature preferences that might differ as a function of random seed, we show that base-model family has systematic, persistent effects. First, we demonstrate that family-level differences appear across minor versions and two orders of magnitude in size. Second, we show persistent difference across preference training across many ablations of data source and quantity, suggesting these representations are deeply rooted and resistant to alignment.
@@
The central idea of inverse reinforcement learning (IRL; Ng and Russell 2000) is to infer a reward model from observed behavior, under the assumption that the observed agent is maximizing this reward. In the context of finetuning LLMs with a KL-regularized reward function, a bandit formulation of IRL has a closed-form solution: the key insight behind DPO (Rafailov et al., 2023), which represents the reward model via a parametric policy, allowing one to finetune via supervised learning. The full IRL setting has been derived in Rafailov et al. (2024). Such implicit rewards have been used as targets for reward distillation as part of finetuning algorithms (Gao et al., 2024; Fisch et al., 2024; Nath et al., 2024; Chen et al., 2025a). To the best of our knowledge, we do not know of previous work systematically analyzing the properties of an implicit reward model defined by two pre-existing LLMs.
@@ 6 Limitations & Conclusion
Exhaustive search surfaces provably optimal/pessimal responses within a given length and avoids the need for sampling (and choice of temperature and sampling algorithm) as in more generative forms of evaluation, however short responses restrict the scope of prompts that can be studied, and a focus on the first output token precludes inference-time compute. Token-level analysis also requires care when comparing across tokenizers. However, results generalize across prompt perturbations: we expand the 3 prompts used by Christian et al. (2025) to 54 prompts, and find that the effects we observe generalize across prompt variations. Second, multi-token responses replicate single-token results. Christian et al. (2025), who introduced exhaustive token search, used techniques from the jailbreaking community such as Greedy Coordinate Gradient (GCG) to derive near-optimal model responses at greater lengths (2-token, 9-token, etc.). These reproduce the same qualitative patterns observed in the provably optimal/pessimal single-token responses, offering preliminary evidence that single-token findings generalize to longer sequences.
@@
While our in-house RM training focused on 2B and 3B models (varying data source and quantity rather than model size), our RewardBench results show that the agency/communion difference between Llama and Gemma RMs is observable at sizes ranging from 2B to 27B, and our implicit RM analysis shows robust model-family differences from 1B to 70B, which appear to increase with model size. Deriving formal scaling laws for both model size and data quantity is a key direction for future work. Second, we focus on Llama and Gemma RMs specifically, owing to their prevalence on RewardBench, however our supplementary analysis (Appendix H) extends these findings to Qwen RMs, which exhibit a communion bias even stronger than that of Gemma. An exhaustive survey of open-weight base models, mapping their differences, would be highly valuable. Third, we focus on the moral “Big Two” of agency/communion, though Section 2 shows similar biases in the five dimensions of the MFD2. Future work extending to yet other dimensions of value would enrich the picture. Finally, mechanistic interpretability tools are needed to reveal how exactly values are inherited from pretraining.
@@
Our results pose significant questions for standard alignment practice. While RLHF and related techniques effectively address style, tone, and avoidance of harmful content, the vast quantities of pre-training data – outstripping preference data by many orders of magnitude – create persistent value biases that cannot be readily overcome via preference modeling. To our knowledge, this is the first work demonstrating this empirically. These findings have significant implications for pretraining data filtering, which likely shapes models’ moral “intuitions” far more than previously recognized. Our results suggest that sufficient preference data can narrow base-model gaps (Fig. 4b), but how training data composition at different stages of the RLHF pipeline interacts with pretraining biases remains underexplored. Developing targeted mitigation strategies – including data filtering, reweighting, augmentation, and debiasing – represents vital future work.
@@ Acknowledgments
BC is supported by the Clarendon Fund. JT is supported by a postdoctoral fellowship from the Natural Sciences and Engineering Research Council of Canada (NSERC) [PDF-578249]. HK is supported by the Economic and Social Research Council [ES/P000649/1]. CS is supported by a Wellcome Trust Discovery Award [227928/Z/23/Z] and an ATRAE Award from the Agencia Estatal de Investigación (AEI).
@@
To ensure full reproducibility of our results, all code is available on GitHub at https://github.com/brchristian/reward_models_inherit_value_biases_from_pretraining. This includes training scripts for RM training, prompt generation, exhaustive-token-search inference, computing MWLR scores, and reproducing all figures and statistical tests.
Model checkpoints from our controlled RM training experiments (Section 4) are available on Hugging Face Hub at https://huggingface.co/collections/Oxford-HIPlab/reward-models-inherit-value-biases-from-pretraining-iclr2026.
@@ Appendix B Psycholinguistic Approach: Big Two and MFD2
To quantify the value biases of RMs, and the relevant pretrained LLMs, we borrowed approaches from a branch of psycholinguistics that quantifies the words people use to shed light on their psychological functioning and individual differences (Pennebaker et al., 2003). One prominent computational approach for this relies on counting and statistically analyzing different features of language, using specially compiled corpora (or dictionaries) that code different words for features of interest. These corpora are hand-crafted by human experts and carefully validated through, for instance, investigations of how conclusions drawn from them relate to other behavioral or self-report measures (i.e., does the result of corpus-based analysis agree with the results of a psychological experiment or with participants’ description of themselves?). Here, we focus our analyses on two relevant psycholinguistic corpora that enumerate words relating to several dimensions of human values: the Big Two (Abele and Wojciszke, 2018) and Moral Foundations Theory (Graham et al., 2009).
@@
The Big Two has a rich history in psychology, influencing empirical work and theories of personality, motivation and social functioning (Abele and Wojciszke, 2018). It comprises the constructs “agency” and “communion,” that relate to “fundamental modalities in the existence of living forms, agency for the existence of an organism as an individual, and communion for the participation of the individual in some larger organism of which the individual is part” (Bakan, 1966, pp. 14–15). And so, the terms agency and communion encompass concerns, motivations or values relating to the achievement of individual goals (e.g., freedom, success, ability) and to relationships with others (e.g., love, support, friendship) respectively. They have previously been related to the basic dimensions, “warmth” and “competence,” according to which people perceive, interpret and stereotype social others (Fiske, 2018). The Big Two dictionary was developed and validated by Pietraszkiewicz et al. (2019) to quantify agentic and communal content in natural language, building on seminal work in psychology that has demonstrated gender biases in recommendation letters (Madera et al., 2009), with female candidates being described as more communal and less agentic than their male counterparts.
@@
The Big Two dictionary contains various word fragments with wildcard character (*), representing the potential addition of zero or more additional characters. For instance, achiev* (agency) could denote achieve, achiever, achievement, etc. For the purposes of our analyses, we handcrafted a corpus of plausible completions. We chose to do this, instead of, for instance exhaustively searching for any possible word completions or inflections / “legal” completions to word roots, as those two approaches led to too many degenerate cases (e.g. winter and wing for win*, or compass along with compassionate). This produced an “unrolled” list of 963 words, 162 of which were nouns. We used the full list for our exhaustive token search analyses (on Christian et al. (2025)’s existing RM data and the data from our own RM training) and the list of nouns for the analyses of the 10 RewardBench RMs across 54 prompts in section 2 and the base-model log probabilities in section 3. Our choices here were motivated by several concerns: (1) RMs exhibit relatively lower sensitivity to the grammatical correctness and stylistic variations of prompt responses relative to LLMs (Christian et al., 2025), (leading us to prefer the noun set for the logprob analyses), and (2) RM token evaluation is more computationally expensive, because each token needs to be evaluated in a separate forward pass, (leading us to generally prefer the smaller noun set, unless exhaustive token data was needed for additional analyses).
@@
The Moral Foundational Dictionary (MFD) was originally developed by Graham et al. (2009) to quantify the moral frames and intuitions used in moral texts (e.g., sermon speeches) by conservative vs. liberal public leaders. It comprises a list of words, hand-coded by expert moral psychologists to reflect five moral “intuitions”: harm/care, fairness/reciprocity, ingroup/loyalty, authority/respect, and purity/sanctity. It was subsequently extended and psychometrically validated as the Moral Foundations Dictionary 2 (MFD2) in a replication study by Frimer (2020). Whilst MFD2 codes for both “virtue” and “vice” words along the five moral foundations (i.e., in the case of the authority foundation, “virtue” words track authority, “vice” track subversion), we focused our analyses on “virtue” for tractability.
@@
Appendix C Value preferences from 10 leading RMs based on Gemma and Llama: Big Two and MFD-2

### v2

@@ 4 Dynamics of Inherited Values Over the Course of RM Training
So far, we have shown that existing open-source RMs based on Llama and Gemma exhibit stereotyped value biases for agency and communion (respectively) that can be traced back to the log probabilities of the instruction-tuned and pretrained versions of the base models, as well as represented by the reward scores of the implicit RM they define. To understand how these value biases evolve over the course of RM training, we perform a set of controlled experiments, training our own RMs from different base models while holding all training parameters identical and controlling for various sources and quantities of training data.
@@ 4.1 Experimental Setup
In order to ensure the inheritability of values is not particular to the preference dataset used for training, we train sets of Llama- and Gemma-based RMs using either of two non-overlapping datasets: Skywork v0.2 ($\approx$77k preferences) and Unified Feedback ($\approx$850k preferences). To establish whether more preference data attenuates the inherited value biases from pretraining, we run experiments with various ablations of the Unified Feedback dataset: 13k, 27k, 53k, or 106k. We train Skywork RMs using the full set of 77k preferences.
@@
RMs are initialized either from Llama 3.2 3B Instruct (“Llama”) or Gemma 2 IT 2B (“Gemma”). We train all RMs with identical hyperparameters: 2 epochs using low-rank adaptation (LoRA, Hu et al. 2022) ($\text{rank}=32$, $\alpha=64$) and AdamW optimizer with learning rate 1e-5, effective batch size 16 (minibatch size 4 $\times$ 4 gradient accumulation steps), and maximum sequence length of 1024 tokens, using Bradley-Terry loss. We run with fixed random seeds to ensure reproducibility.
@@ 4.2 Results
We compare the ranked reward scores assigned by Llama- and Gemma-based RMs to agency- and communion-related tokens in the Big Two corpus; that is, this analysis focuses on the intersection of Gemma and Llama token vocabularies and the Big Two corpus. In Fig. 4(a), we plot the evolution of Big Two ranks for the prompt “What, in one word, is the greatest thing ever?” over the course of training with Skywork. First, consistent with the results so far, the Llama RM ranks agency terms higher than its Gemma counterpart, and the Gemma RM ranks communion terms higher than the Llama one. Second, the gap between Gemma and Llama is widest at the start of training and gradually narrows over the first 4 checkpoints. Third, and crucially, this gap does not close: ranks for agency and communion stabilize for both base models about a third of the way through training (see Appendix G for Kendall $\tau$ results).
Refer to caption(a)
Refer to caption(b)
Figure 4: (a) A pair of Llama and Gemma RMs trained using Skywork 80k preference data, checkpointed every 1000 steps during training, evaluated with the prompt “What, in one word, is the greatest thing ever?” (b) Ablation studies for data source (Unified Feedback
$\circ$
vs. Skywork $\triangle$) and quantity (13k, 27k, 53k, 77k, and 106k), depicting final checkpoints of all runs. We show the gap in preference over the Big Two between Llama (blue) and Gemma (red) at the end of training. For comparability, we also include data from Gemma- and Llama-based “GRMs” trained by Yang et al. (2024) using a combination of regularized BT on a 632k mixture of open-source datasets ($\diamond$) plus standard BT on Skywork.
@@
To zoom in on the relative changes during RM training, we compare which tokens change most in reward-score rankings between early (1000) and late (9578) training checkpoints. Based on our previous findings, we would expect that Llama and Gemma RMs inherit initial biases toward agency and communion tokens (respectively), which fade in influence during training, as the two models move closer together. This is exactly what we find (Fig. LABEL:fig:rank_change_first_and_final). Over the course of training, the Gemma RM comes to increase the reward scores it assigns to agency terms like “choice” while decreasing scores for communion terms like “neighbors,” “teachers,” or “volunteers.” Meanwhile, the Llama RM comes to more highly reward communion terms like “compromises,” “marriages,” and “families,” while lowering its scores for agency terms like “accuracy” and “decision.” (Fig. A7 depicts the training dynamics of these tokens.)
@@
Our ablation studies address how the gap between RM ranks for Big Two terms changes across fully trained RMs as a function of data source and quantity. In Fig. 4(b), each dot represents a model at the end of training on a given source and amount of data. Data source does not make a big difference, but additional preference data helps mitigate the bias from pretraining. Approximately 100k or more preference pairs appear necessary to mitigate the difference between Gemma and Llama bases in our experiments. While these findings demonstrate that some base-model biases may be overcome with sufficient quantities of preference data, two caveats are appropriate. First, here we tested two dimensions of value exclusively (from potentially many value dimensions that can be affected by pretraining biases). Even more data may be needed to attenuate pretraining bias in a multi-dimensional value space. Second, here we tested only two specific base models. In fact, in an exploratory extension to our RM training experiments in Appendix H with Qwen-based RMs, we found that even after training on 100k preferences, the gap in relative agency/communion preference between Qwen and either Gemma or Llama RMs does not close.
Finally, even with very large quantities of preference data, the base model can leave a substantial impact. While our in-house RMs were trained with standard Bradley-Terry loss, in Fig. 4(b) we also plot data from Gemma- and Llama-based “Generalizable Reward Models” (GRMs) trained by Yang et al. (2024). Because they keep the base model’s language head and apply a regularizer that preserves the generative capability of the model’s hidden states, it is conceivable that the base-model biases persist more strongly: we see a striking agency/communion gap even after training on more than 630k preferences. More targeted experiments would be needed to understand the interaction of base-model bias and GRM regularization specifically, but this underscores the importance of carefully considering methodological choices when building RMs.
@@ 5 Related Work
Biases from Pretraining
Recent work has highlighted the importance of pretraining for alignment. Maini et al. (2025) show that safeguards during pretraining reduce vulnerability to malicious attacks relative to post-training approaches; they argue post-training requires the model to (ineffectively) “unlearn” harmful patterns acquired in pretraining. O’Brien et al. (2025) and Chen et al. (2025b) demonstrate that filtering pretraining data is effective in reducing risks from adversarial attacks. Qi et al. (2024) argue that current safety finetuning practices are “shallow” and leave models vulnerable to jailbreaks. Korbak et al. (2023) pretrain LLMs in line with human preferences, and demonstrate that this outperforms post-training alignment. These empirical results relate to a stream of research that has demonstrated that models trained with stochastic gradient descent exhibit robust “simplicity biases” (Jain et al., 2024; Shah et al., 2020; Nakkiran et al., 2019), whereby they first learn simpler functions that can explain patterns in the data; exclusion (or over-inclusion) of certain perspectives in pretraining can lead to class imbalances that cause robust biases downstream. Fulay et al. (2024) observe political bias in RMs but leave the source as an open question; Xiao et al. (2025) show that bias can propagate through KL-regularization during post-training and propose mitigations. Our work identifies pretraining as the source of RM bias, and reveals that regularization addresses only half the problem, since RMs themselves inherit biases that directly inform the post-training reward. In concurrent work, Murthy et al. (2026) trace value trade-offs in LLMs through RLHF, and similarly find that the choice of base model has a persistent impact, with the largest gaps occurring at initialization and narrowing, but not converging, over the course of training – just as we find in our own experiments training RMs.
@@
Black et al. (2022) coined the term “model multiplicity” to describe a common phenomenon whereby models perform similarly on a given performance metric while differing significantly in their internal representations or point-wise behavior. Base-model differences despite similar performance on RewardBench fit our work into this literature; however, our findings go beyond typical model multiplicity in important ways. Unlike idiosyncratic feature preferences that might differ as a function of random seed, we show that base-model family has systematic, persistent effects. First, we demonstrate that family-level differences appear across minor versions and two orders of magnitude in size. Second, we show persistent differences through preference training across many ablations of data source and quantity, suggesting these representations are deeply rooted and resistant to alignment.
@@
The central idea of inverse reinforcement learning (IRL; Ng and Russell 2000) is to infer a reward model from observed behavior, under the assumption that the observed agent is maximizing this reward. In the context of finetuning LLMs with a KL-regularized reward function, a bandit formulation of IRL has a closed-form solution: the key insight behind DPO (Rafailov et al., 2023), which represents the reward model via a parametric policy, allowing one to finetune via supervised learning. The full IRL setting has been derived in Rafailov et al. (2024). Such implicit rewards have been used as targets for reward distillation as part of finetuning algorithms (Gao et al., 2024; Fisch et al., 2024; Nath et al., 2024; Chen et al., 2025a). To the best of our knowledge, no previous work has systematically analyzed the properties of an implicit reward model defined by two pre-existing LLMs.
@@ 6 Limitations & Conclusion
Exhaustive search surfaces provably optimal/pessimal responses within a given length and avoids the need for sampling (and choice of temperature and sampling algorithm) as in more generative forms of evaluation; however, short responses restrict the scope of prompts that can be studied. Token-level analysis also requires care when comparing across tokenizers. Nevertheless, results generalize across prompt perturbations: we expand the 3 prompts used by Christian et al. (2025) to 54 prompts with consistent results. Moreover, multi-token responses replicate single-token results. Christian et al. (2025), who introduced exhaustive token search, used techniques from the jailbreaking community such as Greedy Coordinate Gradient (GCG) to derive near-optimal model responses at greater lengths (2-token, 9-token, etc.). These reproduce the same qualitative patterns observed in the provably optimal/pessimal single-token responses, offering preliminary evidence that single-token findings generalize to longer sequences.
@@
While our in-house RM training focused on 2B and 3B models (varying data source and quantity rather than model size), our RewardBench results show that the agency/communion difference between Llama and Gemma RMs is observable at sizes ranging from 2B to 27B, and our implicit RM analysis shows robust model-family differences from 1B to 70B, which appear to increase with model size. Deriving formal scaling laws for both model size and data quantity is a key direction for future work.
We focus on Llama and Gemma RMs specifically, owing to their prevalence on RewardBench, but our supplementary analysis (Appendix H) extends these findings to Qwen RMs, which exhibit a communion bias even stronger than that of Gemma. An exhaustive survey of open-weight base models, mapping their differences, would be highly valuable. Likewise, we focus on the moral “Big Two” of agency/communion, though Section 2 shows similar biases in the five dimensions of the MFD2. Future work extending to yet other dimensions of value would enrich the picture. Finally, mechanistic interpretability tools are needed to reveal how exactly values are inherited from pretraining.
@@
Our results pose significant questions for standard alignment practice. While RLHF and related techniques effectively address style, tone, and avoidance of harmful content, the vast quantities of pretraining data – outstripping preference data by many orders of magnitude – create persistent value biases that cannot be readily overcome via preference modeling. To our knowledge, this is the first work demonstrating this empirically. These findings have significant implications for pretraining data filtering, which likely shapes models’ moral “intuitions” far more than previously recognized. Our results suggest that sufficient preference data can narrow base-model gaps (Fig. 4(b)), but how training data composition at different stages of the RLHF pipeline interacts with pretraining biases remains underexplored. Developing targeted mitigation strategies – including data filtering, reweighting, augmentation, and debiasing – represents vital future work.
@@ Acknowledgments
Thank you to Rui Yang, Owain Evans, Carroll Wainwright, Amanda Askell, and Jordan Fisher for helpful discussions. BC is supported by the Clarendon Fund. JT is supported by a postdoctoral fellowship from the Natural Sciences and Engineering Research Council of Canada (NSERC) [PDF-578249]. HK is supported by the Economic and Social Research Council [ES/P000649/1]. CS is supported by a Wellcome Trust Discovery Award [227928/Z/23/Z] and an ATRAE Award from the Agencia Estatal de Investigación (AEI).
@@
To ensure full reproducibility of our results, all code and data are publicly available. Code for prompt generation, exhaustive-token-search inference, computing MWLR scores, and reproducing all figures and statistical tests is available at https://github.com/brchristian/reward_models_inherit_value_biases_from_pretraining.
Model checkpoints from our controlled RM training experiments (Section 4) are available on Hugging Face Hub at https://huggingface.co/collections/Oxford-HIPlab/reward-models-inherit-value-biases-from-pretraining-iclr2026. Code for RM training is available at https://github.com/brchristian/Generalizable-Reward-Model.
@@ Appendix B Psycholinguistic Approach: Big Two and MFD2
To quantify the value biases of RMs, and the relevant pretrained LLMs, we borrowed approaches from a branch of psycholinguistics that quantifies the words people use to shed light on their psychological functioning and individual differences (Pennebaker et al., 2003). One prominent computational approach for this relies on counting and statistically analyzing different features of language, using specially compiled corpora (or dictionaries) that code different words for features of interest. These corpora are hand-crafted by human experts and carefully validated through, for instance, investigations of how conclusions drawn from them relate to other behavioral or self-report measures (i.e., does the result of corpus-based analysis agree with the results of a psychological experiment or with participants’ descriptions of themselves?). Here, we focus our analyses on two relevant psycholinguistic corpora that enumerate words relating to several dimensions of human values: the Big Two (Abele and Wojciszke, 2018) and Moral Foundations Theory (Graham et al., 2009).
@@
The Big Two has a rich history in psychology, influencing empirical work and theories of personality, motivation and social functioning (Abele and Wojciszke, 2018). It comprises the constructs “agency” and “communion,” that relate to “fundamental modalities in the existence of living forms, agency for the existence of an organism as an individual, and communion for the participation of the individual in some larger organism of which the individual is part” (Bakan, 1966, pp. 14–15). And so, the terms agency and communion encompass concerns, motivations or values relating to the achievement of individual goals (e.g., freedom, success, ability) and to relationships with others (e.g., love, support, friendship), respectively. They have previously been related to the basic dimensions, “warmth” and “competence,” according to which people perceive, interpret and stereotype social others (Fiske, 2018). The Big Two dictionary was developed and validated by Pietraszkiewicz et al. (2019) to quantify agentic and communal content in natural language, building on seminal work in psychology that has demonstrated gender biases in recommendation letters (Madera et al., 2009), with female candidates being described as more communal and less agentic than their male counterparts.
@@
The Big Two dictionary contains various word fragments with wildcard character (*), representing the potential addition of zero or more additional characters. For instance, achiev* (agency) could denote achieve, achiever, achievement, etc. For the purposes of our analyses, we handcrafted a corpus of plausible completions. We chose to do this, instead of, for instance, exhaustively searching for any possible word completions or inflections/“legal” completions of word roots, as those two approaches led to too many degenerate cases (e.g., winter and wing for win*, or compass along with compassionate). This produced an “unrolled” list of 963 words, 162 of which were nouns. We used the full list for our exhaustive token search analyses (on Christian et al. (2025)’s existing RM data and the data from our own RM training) and the list of nouns for the analyses of the 10 RewardBench RMs across 54 prompts in Section 2 and the base-model log probabilities in Section 3. Our choices here were motivated by several concerns: (1) RMs exhibit lower sensitivity than LLMs to the grammatical correctness and stylistic variations of responses (Christian et al., 2025) (leading us to prefer the noun set for the logprob analyses), and (2) RM token evaluation is more computationally expensive, because each token must be evaluated in a separate forward pass (leading us to generally prefer the smaller noun set unless exhaustive token data were needed for additional analyses).
@@
The Moral Foundations Dictionary (MFD) was originally developed by Graham et al. (2009) to quantify the moral frames and intuitions used in moral texts (e.g., sermon speeches) by conservative vs. liberal public leaders. It comprises a list of words, hand-coded by expert moral psychologists to reflect five moral “intuitions”: harm/care, fairness/reciprocity, ingroup/loyalty, authority/respect, and purity/sanctity. It was subsequently extended and psychometrically validated as the Moral Foundations Dictionary 2 (MFD2) in a replication study by Frimer (2020). While MFD2 codes for both “virtue” and “vice” words along the five moral foundations (i.e., in the case of the authority foundation, “virtue” words track authority, “vice” track subversion), we focused our analyses on “virtue” for tractability.
@@
Appendix C Value Preferences from 10 Leading RMs Based on Gemma and Llama: Big Two and MFD2